RC Meeting, Tuesday, October 12 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote https://bluejeans.com/861439920

Accelerator Status

  1. Setup for Hall A will started around 4PM yesterday and after some difficulties we received beam back around 9:30 last night.
    • The setup will continue this swing shift when Hall A will lock up, although we should expect beam for most of their setup. We will probably have higher chances of trips due to FSDs originating in the Hall A line.

Hall C Status

  1. Stephen and Dave moved the HMS yesterday to 58.535 degreees (https://logbooks.jlab.org/entry/3922017)
  2. So far we have collected approximately 280 pions at this new kinematic (https://logbooks.jlab.org/entry/3922645)

Current Run Plan

  1. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for magnets: next opportunistic access (Joe Beaufait) ( PENDING )
  2. The HB has seen some fluctuations in the LHe levels. At the next opportunistic access she would like to investigate the cryo line. ( PENDING )
  3. Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
  4. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.

Special Requests

Useful Links